OpenStudy (anonymous):

Use Euler's Formula to find the missing number. Faces:_____ Edges:15 Vertices: 9

OpenStudy (zzr0ck3r):

For a planar graph we have \(v-e+f=2\) So you have \(9-15+f=2\)

OpenStudy (zzr0ck3r):

Solve for \(f\).

OpenStudy (anonymous):

8! thank you can you help me with another please?

OpenStudy (anonymous):

Use Euler's Formula to find the missing number. Faces:8 Edges:____ Vertices: 6 @zzr0ck3r

OpenStudy (anonymous):

@jim_thompson5910 @izuru @DanJS

OpenStudy (anonymous):

@foreverthebeast @FourJustice @animalfan1 @LynFran @nuttyliaczar

OpenStudy (zzr0ck3r):

Just plug them in....

OpenStudy (zzr0ck3r):

Plug the ones you are given into \(v-e+f=2\). Solve for the only variable left.

OpenStudy (anonymous):

LAst one please and thank you for the help

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

Just imagine it. What do you think the hole in the paper will be?